# White Men Lead, Black Women Help? Benchmarking Language Agency Social Biases in LLMs

### v4

@@ 4.1 Findings 1: LLM generations are More Gender Biased than Human-Written Texts
We establish comparison with bias in LLM-generated texts by incorporating analysis on 3 existing datasets: human-written biographies in Bias in Bios, human-written professor reviews on RateMyProfessor, and the reference letter dataset in  Wan et al. (2023)’s work, which consists of letters generated by LLMs given extensive biographical information (e.g. multi-sentence descriptions of career development) about specific individuals. Since we do not find any publicly available large-scale dataset for reference letters,  Wan et al. (2023)’s data is our best choice as a proxy of human-written letters. Additionally, no openly-accessible datasets with racial information were found in our search, limiting our analysis to gender biases.
@@ 4.1.1 Human-Written Texts: Dataset Details
Reference Letters   Since we were not able to find publicly available human-written reference letter datasets, we choose to use the reference letter dataset from the Context-Based Generation (CBG) setting in  Wan et al. (2023)’s work. The CBG setting provides a paragraph of biographical information about individuals (e.g. career, life) to prompt LLMs for letter generations, which is very similar to real-world reference-letter-writing scenarios. Therefore, we use  Wan et al. (2023)’s dataset as a proxy for human-written reference letters.
Figure 3: Visualization of the average ratio gap between agentic and communal sentences in the 3 datasets for different intersectional genders and racial groups.
@@ 4.1.2 Comparison Results
Figure 2 visualizes language agency gender biases in human-written and LLM-generated biographies, professor reviews, and reference letters. We report the gender differences (Male - Female) in the intra-group agency-communal ratio gaps. Quantitative results are in Appendix D, Table 14. Below are our observations:
@@
Biases observed in human-written texts in our study align with findings of social science studies. We stratify analysis on the human-written biography dataset based on professions in Appendix D, and found that occupations with greatest biases—such as pastor, architect, and software engineer—are also reported by real-world studies to be male-dominated (Kathleen Schubring, ; A. Nicholson et al., ; Kaminski, ). Academic departments in which the highest language agency biases in professor reviews are identified—such as Accounting, Sociology, and Chemistry—have also been proven for male dominance (200, 2009; Girgus, ; Seijo, ). Alignment between our observations and real-world inequalities further shows the effectiveness of agency in capturing social biases.
@@ 4.2 Findings 2: LLMs Suffer From Gender, Racial, and Especially Intersectional Biases in Language Agency
Model Text Type Problem Type Race Gender Diff. (M-F) Gender Diff. (post-mitigate)
ChatGPT Biography Amplification Black 3.38 18.68
Professor Review Overshooting White 5.56 -16.84
Black 2.41 -9.91
Mistral Professor Review Amplification White 7.02 14.59
Black 8.02 17.93
Asian 7.99 23.00
Reference Letter Amplification Black 4.42 12.95
Asian 10.87 15.92
Llama3 Biography Amplification White 6.06 10.59
Black 6.21 11.24
Table 2: Selected quantitative results showing problems with bias amplification and bias overshooting for prompt-based mitigation methods in LLMs. For each racial group, we report the gender difference (male-female) in agency-communal ratio gap in texts. ChatGPT demonstrates both amplification and overshooting, whereas Mistral and Llama mainly suffer from amplified language agency bias levels.
@@
In intersectional bias dimension, texts depicting individuals at the intersection of gender and racial minority groups—such as Black females—possess remarkably lower language agency levels. Both quantitative results in Appendix D Tables 21, 23,25 and visualized illustrations in Figure  3 show severe intersectional biases across all LLMs on all generation tasks—those who are at the intersection of gender and racial minority groups are the most vulnerable to biases in language agency. For instance, ChatGPT- and Llama3-generated reviews for Black female professors show the lowest level of agency across all intersectional groups. Interestingly, we observe that on all text generation tasks, language agency is notably higher in texts about males within each racial group (e.g. Black males are described with more agentic language than Black females). These observations further align with prior social science findings on intersectional biases targeting gender and racial minority groups in texts (Field et al., 2022; Adams et al., 2019; Lemieux et al., 2023; Otterbacher, 2015; Chávez and Mitchell, 2020).
4.3 Findings 3: All LLMs Show Agency Biases, and Llama3 is the Most Biased One
As shown by the overall quantitative results (bottom right cell in rows for each model) of biases in each dimension, for each text generation task, and for all 3 models, Llama3 suffers from the most remarkable language agency biases. Through inspecting individual bias categories, it is shocking to see that Llama3 in fact carries the highest level of biases in 2 of the 3 investigated text generation tasks—biography and reference letter—across gender, race, and intersectional dimensions. This warns that although recently-developed LLMs all demonstrate fascinating text generation abilities, they can demonstrate drastically different levels of fairness issues. Using such technology without scrutiny can result in the propagation of severe social harm.
4.4 Findings 4: Simple Prompt-Based Mitigation Could Worsen Language Agency Bias.
@@ 4.3 Findings 3: All LLMs Show Agency Biases, and Llama3 is the Most Biased One
A large body of recent research has explored the use of “ethical intervention”, or prompt-based mitigation, to resolve fairness issues in textual and multimodal generative models (Bansal et al., 2022; Ganguli et al., 2023; Huang et al., 2024; Wan and Chang, 2024). Following the previous studies, we experimented with a prompt-based bias mitigation method by appending a “fairness instruction” at the end of each text generation prompt: “When generating the {text type}, ensure that you display no biases in language agency across gender or race.”
Quantitative results in Table 1 show that prompt-based methods fail to effectively resolve language agency bias. More shockingly, we reveal that fairness instructions in prompts could even result in higher bias levels in LLM-generated texts. We found 2 main problems with prompt-based mitigation methods in LLMs: (1) bias amplification, for which even more severe biases are observed, and (2) bias overshooting, where biases are shifted towards anti-stereotype directions (e.g. biased towards males). In Table  2, we provide quantitative examples for these problems for the 3 LLMs.
These surprising findings unveil the severe drawbacks of prompt engineering as a bias mitigation method—LLMs might fail to understand what is “fair” in language agency, therefore worsening existing biases or resulting in biases in the anti-stereotypical direction.
@@ 5 Conclusion
In this work, we propose the Language Agency Bias Evaluation (LABE) framework to systematically and comprehensively measure gender, racial, and intersectional biases in language agency across a wide scope of text generation tasks. To build better agency evaluation tools, we also contribute the Language Agency Classification (LAC) dataset for training accurate language agency classifiers. Through experimenting on 3 LLMs, we found that: (1) LLM-generated texts often carry remarkably higher levels of bias than human-written language; (2) People who are at the intersection of gender and racial minority groups (e.g. Black females) are the most vulnerable to language agency biases; (3) Compared with ChatGPT and Mistral, Llama3’s outputs tend to show the greatest overall level of language agency bias; (4) Simple prompt-based mitigation methods might result in the amplification and overshooting of biases, worsening the fairness issue in LLMs. Our LABE benchmark addresses previous research gaps and provides valuable technical contributions, and our findings point towards the imminence of comprehensively examining and resolving language agency biases in LLMs, forewarning potential social risks of using LLM-generated texts without scrutiny.
@@
Experiments in this study incorporate language models that were pre-trained on a wide range of text from the internet and have been shown to learn or amplify biases from the data used. Since we utilize a language model to synthesize a language agency classification dataset, we adopt a number of methods to prevent potential harm and bias propagation: (1) we prompt the model to paraphrase each input into an agentic version and a communal version, ensuring the balance in the preliminary generated dataset, and (2) we invite expert annotators to re-annotate the generated data, to verify and ensure the quality of the final dataset used to train language agency classifiers. We encourage future extensions of our works to also consider this factor in their research, so as to draw reliable and trustworthy research conclusions.
@@ Appendix A Prompt Construction for Text Generation Tasks in LABE
We adapt the Context-Less Generation (CLG) approach proposed by  Wan et al. (2023) to prompt LLMs on biography, professor review, and reference letter generation tasks using descriptor-based prompt templates. All prompts contain gender and race descriptors to include demographic information. Additional information such as name, occupation, age, and department are also injected in prompts for different tasks. Table 3 demonstrates prompt templates used in LABE to obtain LLM outputs on the 3 text generation tasks about people with different demographic identities, for comprehensive bias analysis. Table 5 presents lists of descriptors including gender, race, occupations, departments, and ages. To identify natural names for males and females in different racial groups, we prompted ChatGPT to generate $5$ popular male names and $5$ popular female names in each of the $4$ race categories: “White”, “Black”, “Hispanic”, and “Asian”. Table 4 shows the full list of investigated racial groups and corresponding popular names for males and females within each group.
@@ B.6 Model Choices
Generative Model For the Llama2 generative model, we experiment with $4$ different settings: zero-shot prompting without definition, zero-shot prompting with definition, one-shot prompting with definition and an example, and parameter-efficient fine-tuning with LoRA Hu et al. (2021). For reproducibility, we provide the full prompts used to probe Llama2 in zero-shot and few-shot settings in Table 10. For LoRA fine-tuning, we use a learning rate of $5e^{-5}$ to train for $5$ epochs. During inference, we follow the default generation configuration to set top-p to $1.0$, tok-k to $50$, and temperature to $1.0$.
@@
(Wan et al., 2023) 109M MIT License + Fine-Tune $66.49$ $66.49$ $64.22$ $64.82$
@@ B.7 Model Performance
We report the performances of baseline methods to classify language agency, as well as our trained classifiers on the LAC dataset. For baseline methods, we experimented on string matching, sentiment classification, and the agency classifier proposed in Wan et al. (2023)’s work. For string matching, we utilized Stahl et al. (2022)’s released lists of agentic and communal words with no licensing information. For sentiment classification, we utilized the sentiment classification pipeline in the transformers library with the off-the-shelf “distilbert/distilbert-base-uncased-finetuned-sst-2-english” [^3] https://huggingface.co/distilbert/distilbert-base-uncased-finetuned-sst-2-english model.
@@ Appendix C Human-Written Datasets Details
Reference Letter (Wan et al., 2023) We directly utilize the reference letter dataset generated under the CBG setting in  Wan et al. (2023)’s work, which is released under the MIT License.
Appendix D Additional Experiment Results
@@
D.1 Quantitative Results: Bias in LLMs vs. Human-Written Texts
@@
D.2 Stratified Analysis on Human-Written Datasets
We stratify analysis on the human-written biography dataset based on professions and provide full results in Table 16. We then visualize the top $8$ most biased occupations as overlap horizontal bar graphs in Figure 4. Drastic language agency gender biases are found for pastor, architect, and software engineer. Interestingly, real-world reports have also demonstrated male dominance and gender bias in these occupations (Kathleen Schubring, ; A. Nicholson et al., ; Kaminski, ). Similarly, we stratify our analysis on the human-written professor review dataset based on academic departments in Table 15, and visualize the top $8$ most biased departments in Figure 4. Greatest biases are observed in reviews for professors in departments such as Accounting, Sociology, and Chemistry; all $3$ departments have been proven to be male-dominated (200, 2009; Girgus, ; Seijo, ). Language agency gender biases found on human-written texts in our study align with findings of social science studies, showing that our proposed evaluation tools effectively capture implicit language style biases.

### v5

@@ 4.1 Findings 1: LLM generations are More Gender Biased than Human-Written Texts
We establish comparison with bias in LLM-generated texts by incorporating analysis on 3 existing datasets: human-written biographies in Bias in Bios, human-written professor reviews on RateMyProfessor, and the reference letter dataset in  Wan et al. (2023a)’s work, which consists of letters generated by LLMs given extensive biographical information (e.g. multi-sentence descriptions of career development) about specific individuals. Since we do not find any publicly available large-scale dataset for reference letters,  Wan et al. (2023a)’s data is our best choice as a proxy of human-written letters. Additionally, no openly-accessible datasets with racial information were found in our search, limiting our analysis to gender biases.
@@ 4.1.1 Human-Written Texts: Dataset Details
Reference Letters   Since we were not able to find publicly available human-written reference letter datasets, we choose to use the reference letter dataset from the Context-Based Generation (CBG) setting in  Wan et al. (2023a)’s work. The CBG setting provides a paragraph of biographical information about individuals (e.g. career, life) to prompt LLMs for letter generations, which is very similar to real-world reference-letter-writing scenarios. Therefore, we use  Wan et al. (2023a)’s dataset as a proxy for human-written reference letters.
Figure 3: Visualization of the average ratio gap between agentic and communal sentences in the 3 datasets for different intersectional genders and racial groups. We observe that across generation tasks, texts generated for minority gender, racial, and intersectional groups tend to demonstrate low agency levels. For instance, Black female professors receive reviews with the lowest agency for both ChatGPT- and Llama3-generated reviews.
@@ 4.1.2 Comparison Results
Figure 2 visualizes language agency gender biases in human-written and LLM-generated biographies, professor reviews, and reference letters. We report the gender differences (Male - Female) in the intra-group agency-communal ratio gaps. Quantitative results are in Appendix E.2, Table 16. Below are our observations:
@@
Biases observed in human-written texts in our study align with findings of social science studies. We stratify analysis on the human-written biography dataset based on professions in Appendix E.2, and found that occupations with greatest biases—such as pastor, architect, and software engineer—are also reported by real-world studies to be male-dominated (Kathleen Schubring, ; A. Nicholson et al., ; Kaminski, ). Academic departments in which the highest language agency biases in professor reviews are identified—such as Accounting, Sociology, and Chemistry—have also been proven for male dominance (Kamath et al., 2009; Girgus, ; Seijo, ). Alignment between our observations and real-world inequalities further shows the effectiveness of agency in capturing social biases.
@@ 4.2 Findings 2: LLMs Suffer From Gender, Racial, and Especially Intersectional Biases in Language Agency
In intersectional bias dimension, texts depicting individuals at the intersection of gender and racial minority groups—such as Black females—possess remarkably lower language agency levels. Both quantitative results in Appendix E.2 Tables 23, 25,27 and visualized illustrations in Figure  3 show severe intersectional biases across all LLMs on all generation tasks—those who are at the intersection of gender and racial minority groups are the most vulnerable to biases in language agency. For instance, ChatGPT- and Llama3-generated reviews for Black female professors show the lowest level of agency across all intersectional groups. Interestingly, we observe that on all text generation tasks, language agency is notably higher in texts about males within each racial group (e.g. Black males are described with more agentic language than Black females). These observations further align with prior social science findings on intersectional biases targeting gender and racial minority groups in texts (Field et al., 2022; Adams et al., 2019; Lemieux et al., 2023; Otterbacher, 2015; Chávez and Mitchell, 2020).
Figure 4: Visualization of the proposed Mitigation via Selective Rewrite (MSR) pipeline.
5 Mitigating Language Agency Biases
To investigate whether we can effectively reduce language agency biases, we conducted small-scale experiments with 96 randomly-sampled evaluation prompts for each generation task.
5.1 Prompt-Based Mitigation
@@
Recent research explored the use of “ethical intervention”, or prompt-based mitigation, to resolve fairness issues in textual and multimodal generative models (Bansal et al., 2022; Ganguli et al., 2023; Huang et al., 2024; Wan and Chang, 2024). We experimented with a prompt-based bias mitigation method by appending a “fairness instruction” at the end of each generation prompt.
Quantitative results in Table 2 show that prompt-based methods fail to stably and effectively resolve language agency bias, and could even result in higher bias levels in LLM-generated texts. This shows that simple prompt engineering is not enough as a bias mitigation method—LLMs lack knowledge on how to make fairness improvements, resulting in unstable and suboptimal mitigation results, sometimes even worsening existing biases.
5.2 Mitigation via Selective Rewriting
Model Text Type Bias Dimension
Gender Race Intersect. Overall
ChatGPT Biography 21.93 44.66 130.89 65.83
 +Prompt Mit. 29.55 14.09 45.94 29.86
 +MSR 6.59 26.63 66.72 33.31
Professor Review 6.35 36.66 86.83 43.28
 +Prompt Mit. 15.50 34.90 62.26 37.55
 +MSR 9.91 18.58 29.42 19.30
Reference Letter 40.99 13.57 43.65 32.74
 +Prompt Mit. 3.15 51.36 62.79 39.10
 +MSR 13.33 15.95 29.49 19.59
Mistral Biography 45.08 9.05 73.09 42.41
 +Prompt Mit. 29.22 28.19 43.59 33.67
 +MSR 17.40 7.72 53.27 26.13
Professor Review 0.05 46.22 49.59 31.95
 +Prompt Mit. 103.77 22.85 96.83 74.48
 +MSR 31.84 32.15 127.43 63.81
Reference Letter 38.76 14.96 47.23 33.65
 +Prompt Mit. 89.50 62.29 107.38 86.39
 +MSR 31.16 7.82 29.64 22.87
Llama3 Biography 14.28 34.52 50.72 33.17
 +Prompt Mit. 60.12 80.14 117.50 85.92
 +MSR 18.47 31.42 66.24 38.71
Professor Review 16.26 73.82 90.49 60.19
 +Prompt Mit. 2.85 8.67 16.92 9.48
 +MSR 8.92 73.48 137.46 73.29
Reference Letter 21.60 49.92 61.52 44.35
 +Prompt Mit. 27.58 23.95 49.33 33.62
 +MSR 1.66 8.20 31.14 13.67
Table 2: Experiment results for original LLM-generated texts, prompt-based mitigation, and the proposed MSR method on the 96 sampled evaluation entries. Lowest bias for each task for each LLM is underlined.
Observing the prevalence of language agency bias in LLMs and the unsatisfactory performance of previously introduced prompt-based mitigation methods, we propose Mitigation via Selective Rewriting (MSR), a novel bias mitigation method that utilizes the LAC classifier to identify and revise communal sentences to make them more agentic.
Figure 4 shows the general pipeline of our MSR method. Prompts used for querying revision is provided in Appendix D, Table 15. Using the LAC classifier, MSR identifies communal sentences (as highlighted in red) in LLM-generated texts. Then, we prompt the model to provide a rewrite for the identified parts in the texts to make the language more agentic. The updated generation will then possess a higher overall agency level, addressing the problem of low language agency for minority demographic groups.
Quantitative experiment results in Table  2 prove the effectiveness of MSR in reducing language agency bias compared to the prompt-based method: MSR is able to achieve the lowest overall bias level in 5 out of 9 total task completions across 3 LLMs, whereas prompt-based method only achieves best results in 2 completions. Qualitative examples in Figure 5 show how MSR is able to conduct targeted revisions on LLM-generated texts to only edit communal parts and make them more agentic—for instance, by adding that the professor “implements effective teaching strategies”.
Figure 5: Qualitative results showing the effectiveness of MSR. MSR outperforms prompt-based mitigation by conducting targeted and more controllable edits to make communal parts in the texts more agentic.
Behavioral Analysis
Despite MSR bringing remarkable improvement to bias mitigation, we also observe that neither of the mitigation methods can fully achieve stable and effective bias reduction results—there exist cases where mitigation approaches result in higher bias in generations.
To better understand the behavior and limitations of our MSR mitigation approach, we analyze detailed experiment results on professor reviews generated by Mistral, as provided in Table 3. While our method boosts the average agency levels in professor reviews generated for all social groups, for minority groups like black females, MSR was not able to boost the agency level to as high as that for majority groups such as white males. This indicates that a stronger mitigation method might be needed to effectively remove bias for minority groups. Observations on MSR’s behaviors also explain a rise in variance in agency level across intersectional groups after mitigation, as the boost in agency levels post-mitigation might be more salient for majority groups compared to minority groups. Nevertheless, our results in Table 2 show that MSR is by far among the best mitigation strategies to achieve overall bias removal.
Our detailed observations further highlight the importance of future work to develop new bias mitigation approaches to address the complicated bias in language agency.
Racial Group Gender Setup Agency %
White Male Original 47.86
+ MSR 64.73
Female Original 44.36
+ MSR 61.12
Black Male Original 41.96
+ MSR 68.65
Female Original 37.95
+ MSR 51.34
Hispanic Male Original 40.41
+ MSR 54.66
Female Original 35.64
+ MSR 59.66
Asian Male Original 44.10
+ MSR 62.62
Female Original 40.10
+ MSR 62.64
Table 3: Percentages of agentic sentences for texts generated across racial and gender groups, before and after applying our MSR mitigation method.
@@ 6 Conclusion
In this work, we propose the Language Agency Bias Evaluation (LABE) framework to systematically and comprehensively measure gender, racial, and intersectional biases in language agency across a wide scope of text generation tasks. To build better agency evaluation tools, we also contribute the Language Agency Classification (LAC) dataset for training accurate language agency classifiers. Through experimenting on 3 LLMs, we found that: (1) LLM-generated texts often carry remarkably higher levels of bias than human-written language; (2) People who are at the intersection of gender and racial minority groups (e.g. Black females) are the most vulnerable to language agency biases; (3) Simple prompt-based mitigation methods might result in the amplification and overshooting of biases, worsening the fairness issue in LLMs. Based on empirical observations, we further propose the Mitigation via Selective Rewrite (MSR) method to reduce bias through selectively revising communal parts in model-generated texts to make the language more agentic. Results show the effectiveness of MSR in improving fairness in language agency, but also highlight the importance of future works to develop more controllable and effective bias mitigation approaches.
Acknowledgements
We thank the UCLA-NLP+ members, conference reviewers, and conference chairs for their invaluable feedback. We also hope to acknowledge NSF #2331966 and ONR grant N00014-23-1-2780 for supporting this work.
@@ Acknowledgements
Experiments in this study incorporate Large Language Models that were pre-trained on a wide range of text from the internet and have been shown to learn or amplify biases from the data used (Wan et al., 2023a, b). Since we utilize a language model to synthesize a language agency classification dataset, we adopt a number of methods to prevent potential harm and bias propagation: (1) we prompt the model to paraphrase each input into an agentic version and a communal version, ensuring the balance in the preliminary generated dataset, and (2) we invite expert annotators to re-annotate the generated data, to verify and ensure the quality of the final dataset used to train language agency classifiers. We encourage future extensions of our works to also consider this factor in their research, so as to draw reliable and trustworthy research conclusions.
@@ Appendix A Prompt Construction for Text Generation Tasks in LABE
We adapt the Context-Less Generation (CLG) approach proposed by  Wan et al. (2023a) to prompt LLMs on biography, professor review, and reference letter generation tasks using descriptor-based prompt templates. All prompts contain gender and race descriptors to include demographic information. Additional information such as name, occupation, age, and department are also injected in prompts for different tasks. Table 4 demonstrates prompt templates used in LABE to obtain LLM outputs on the 3 text generation tasks about people with different demographic identities, for comprehensive bias analysis. Table 6 presents lists of descriptors including gender, race, occupations, departments, and ages. To identify natural names for males and females in different racial groups, we prompted ChatGPT to generate $5$ popular male names and $5$ popular female names in each of the $4$ race categories: “White”, “Black”, “Hispanic”, and “Asian”. Table 5 shows the full list of investigated racial groups and corresponding popular names for males and females within each group.
@@ B.6 Model Choices
Generative Model For the Llama2 generative model, we experiment with $4$ different settings: zero-shot prompting without definition, zero-shot prompting with definition, one-shot prompting with definition and an example, and parameter-efficient fine-tuning with LoRA Hu et al. (2021). For reproducibility, we provide the full prompts used to probe Llama2 in zero-shot and few-shot settings in Table 11. For LoRA fine-tuning, we use a learning rate of $5e^{-5}$ to train models for $5$ epochs. During inference, we follow the default generation configuration to set top-p to $1.0$, tok-k to $50$, and temperature to $1.0$.
@@
(Wan et al., 2023a) 109M MIT License + Fine-Tune $66.49$ $66.49$ $64.22$ $64.82$
@@ B.7 Model Performance
We report the performances of baseline methods to classify language agency, as well as our trained classifiers on the LAC dataset. For baseline methods, we experimented on string matching, sentiment classification, and the agency classifier proposed in Wan et al. (2023a)’s work. For string matching, we utilized Stahl et al. (2022)’s released lists of agentic and communal words with no licensing information. For sentiment classification, we utilized the sentiment classification pipeline in the transformers library with the off-the-shelf “distilbert/distilbert-base-uncased-finetuned-sst-2-english” [^3] https://huggingface.co/distilbert/distilbert-base-uncased-finetuned-sst-2-english model.
@@ Appendix C Human-Written Datasets Details
Reference Letter (Wan et al., 2023a) We directly utilize the reference letter dataset generated under the CBG setting in  Wan et al. (2023a)’s work, which is released under the MIT License.
Appendix D Additional Mitigation Details
We hereby provide details on the implementation of mitigation approaches in this study. Table 15 presents the full prompts used for both vanilla prompt-based mitigation and our proposed MSR method.
Prompt-based Mitigation “When generating the {text type}, ensure that you display no biases in language agency across gender or race.”
MSR “Task: Given the following document and a list indicating whether each sentence in the text is agentic (more achievement-oriented, e.g. assertive, leadership) or communal (more social or service-oriented, e.g. friendly, warm), make edits to communal sentences by identifying phrases that demonstrate communal traits and provide rephrases to make each of them more agentic. Example: Text: Jenny excels at her work as a lawyer. She is also a wonderful person to work with, as she is always warm and helpful to all her coworkers. Agency List: [’agentic’, ’communal’]. "edits": ["original":"a wonderful person to work with", "edited":"a trustworthy co-worker","original":"she is always warm and helpful to all her coworkers", "edited":"she always make constructive contributions and demonstrate great leadership in her team"…] Text: (original text). Agency List: (agency list). List of edits (json format):”
Table 15: Prompt for both mitigation methods experimented.
Appendix E Additional Experiment Details
E.1 Additional Implementation Details
Table 10 shows the prompt templates that we use to query the generation of LLMs for different text generation tasks.
E.2 Additional Experiment Results
@@
E.2.1 Quantitative Results: Bias in LLMs vs. Human-Written Texts
@@
E.2.2 Stratified Analysis on Human-Written Datasets
We stratify analysis on the human-written biography dataset based on professions and provide full results in Table 18. We then visualize the top $8$ most biased occupations as overlap horizontal bar graphs in Figure 6. Drastic language agency gender biases are found for pastor, architect, and software engineer. Interestingly, real-world reports have also demonstrated male dominance and gender bias in these occupations (Kathleen Schubring, ; A. Nicholson et al., ; Kaminski, ). Similarly, we stratify our analysis on the human-written professor review dataset based on academic departments in Table 17, and visualize the top $8$ most biased departments in Figure 6. Greatest biases are observed in reviews for professors in departments such as Accounting, Sociology, and Chemistry; all $3$ departments have been proven to be male-dominated (Kamath et al., 2009; Girgus, ; Seijo, ). Language agency gender biases found on human-written texts in our study align with findings of social science studies, showing that our proposed evaluation tools effectively capture implicit language style biases.